# GraphCLIP: Enhancing Transferability in Graph Foundation Models for Text-Attributed Graphs

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 3.1. Graph-Summary Pair Generation
Considering scalability, we employ a sampling function, random walk with restart sampler, to sample subgraphs $\{G_{n}\}_{n=1}^{N}$ from a large graph. These subgraphs will be incorporated into our prompt template and designed instructions to enable LLMs to generate the corresponding graph summaries $\{S_{n}\}_{n=1}^{N}$. In this study, we utilize open LLM, QWen2-72B (Yang et al., 2024a), as the graph summary generator. Combining realistic unlabelled graph data, we generate and curate large-scale graph-summary pair data [^1] https://github.com/ZhuYun97/GraphCLIP/summary/ which contains over 0.2B tokens across academic (He et al., 2023; Wang et al., 2020a; Yang et al., 2016), e-commerce (Hu et al., 2020), and social domains (Huang et al., 2024). For detailed instructions for graph summary generation, please refer to the Appendix C.
@@ 4.1. Datasets
Table 1. Statistics of Text-Attributed Graph datasets. $\mathcal{G}_{\text{source}}$ denotes source datasets, and $\mathcal{G}_{\text{target}}$ indicates target datasets
@@ 4.3.1. Node Classification
For the LLM baselines, we directly apply them to the source data, as they have already been pretrained on extensive corpora, and we find that continuing to pretrain them on our source data deteriorates performance on the target data. For GraphGPT, we utilize their released checkpoint [^2] https://github.com/HKUDS/GraphGPT and conduct experiments under our settings. For other methods, we use their official codes and pretrain models using the source datasets specified in Table 1 before applying them to the target datasets. For data splitting, we use the public split for the WikiCS dataset, while for other datasets, we randomly select 20% as testing samples. We report the mean accuracy along with the standard deviation after five runs with different random seeds.
@@
Table 3. Zero-shot inference for link prediction of different methods across target datasets
@@ Appendix A The Dilemma of Vanilla Contrastive Loss
Consider a binary classification scenario with data $(X_{1},X_{2})\sim\mathcal{N}(0,I_{2})$. When $X_{1}\geq 0$, the label is set to $Y=1$, and the data augmentation process involves multiplying $X_{2}$ by standard normal noise:
(12) $\tau_{\theta}(X)$ $=\left(X_{1},\theta\cdot X_{2}\right)$
$\theta$ $\sim\mathcal{N}(0,1)$
The resulting transformation-induced domain set is defined as =$\mathcal{B}=\left\{\mathcal{G}_{m}:\mathcal{G}_{m}=\left(X_{1},m\cdot X_{2}\right)\text{ for }m\in\mathbb{R}\right\}$. By considering the 0-1 loss, for every $\zeta\geq 0$, there exists a representation $g$ and two domains $\mathcal{G}_{m}$ and $\mathcal{G}_{m^{\prime}}$ such that
@@
yet $g$ performs very differently across the domains $\mathcal{G}_{m}$ and $\mathcal{G}_{m^{\prime}}$:
@@
where $\mathcal{L}_{\text{AL}}$ denotes alignment loss in the vanilla contrastive loss (Wang and Isola, 2020), $\mathcal{R}$ denotes supervised risk of binary classification. This example [^3] For simplicity, we assume the adjacency matrix is an identity matrix in this case. underscores that a representation with a small contrastive loss can still demonstrate significant performance variability across different augmentation-induced domains. The core concept illustrated here is that a low $\mathcal{L}_{\text{AL}}$ is achieved by averaging alignments across different augmentation-induced domains, rather than ensuring uniform alignment. As a result, the representation may experience substantial alignment losses in certain less frequently selected domains.
Proof. For $\zeta\geq 0$, let $t=\sqrt{\zeta}/2$ and $g(x_{1},x_{2})=x1+tx_{2}.$ Then, the alignment loss of $g$ satisfies:
(15) $\mathcal{L}_{\text{AL }}(g;\mathcal{G},\pi)=t^{2}\mathbb{E}X_{2}^{2}\underset{\left(\theta_{1},\theta_{2}\right)\sim\mathcal{N}(0,1)^{2}}{\mathbb{E}}\left(\theta_{1}-\theta_{2}\right)^{2}=2t^{2}<\zeta.$
Set $c$ as 0 and $c^{\prime}$ as $1/t$, it is obviously that:
(16) $\mathcal{R}\left(g;\mathcal{G}_{c}\right)=0$
@@
(17) $\mathcal{R}\left(g;\mathcal{G}_{c^{\prime}}\right)=$
$P\left(X_{1}<0,X_{1}+X_{2}\geq 0\right)+P\left(X_{1}\geq 0,X_{1}+X_{2}\leq 0\right)=\frac{1}{4}$
@@ Appendix B Theoretical Analysis of Invariant Alignment Loss
When $\mathcal{L}_{\text{IAL}}$ is minimized to a small value, it signifies that $\mathcal{R}\left(c\circ g;\mathcal{G}\tau\right)$ remains consistent across various augmentation functions $\tau$, suggesting that the optimal representation for $\mathcal{G}\tau$ closely resembles that of $\mathcal{G}{\tau^{\prime}}$. In other words, representations with smaller $\mathcal{L}_{\text{IAL}}$ are more likely to yield similar linear optimal predictors across different domains, a characteristic that the original alignment loss lacks.
@@ Appendix D Datasets
WikiCS (Mernyei and Cangea, 2020) dataset is a Wikipedia-based dataset designed for benchmarking Graph Neural Networks, comprising 10 computer science branches as classes characterized by high connectivity. Node features are extracted from the corresponding article texts [^4] We obtain the raw texts of each node from https://github.com/pmernyei/wiki-cs-dataset..
@@ G.1. Pretraining Phase
For GraphCLIP, only a few hyperparameters need to be adjusted. In our main experiments, we utilize the AdamW (Loshchilov and Hutter, [n. d.]) optimizer with both learning rate and weight decay set to 1e-5. The graph model employed is GraphGPS (Rampášek et al., 2022), consisting of 12 layers with a hidden size of 1024. For the text model, we use a fine-tuned version of MiniLM  [^5] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 (Wang et al., 2020b), featuring 6 layers with a hidden size of 384. To align both models in a unified subspace, a projector is applied to transform the graph model’s 1024 dimensions to match the 384 dimensions of the text model. During pretraining, we optimize only the parameters of the graph model and projector, keeping the text model frozen to reduce training costs and mitigate catastrophic forgetting. Pretraining is conducted over 30 epochs with a batch size of 800 per GPU, utilizing eight A100-40G GPUs for pretraining within 7 hours. We will release our pretrained checkpoint  [^6] https://github.com/ZhuYun97/GraphCLIP/checkpoints/ after the anonymous phase.
@@ G.2. Zero-shot Learning
In zero-shot learning, we incorporate label information into label-specific sentences to align with the pretraining format. Table 8 presents various prompts designed for different datasets. Placeholders are marked in violet font: “{class}” represents the label text for the target node, and “{class_desc}” is a descriptive sentence generated by LLMs to elaborate on the label. For detailed cases, please refer to our anonymous repository  [^7] https://github.com/ZhuYun97/GraphCLIP/data/load.py.

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/0622footnotetext: Corresponding authors.
@@ 3.1. Graph-Summary Pair Generation
Considering scalability, we employ a sampling function, random walk with restart sampler, to sample subgraphs $\{G_{n}\}_{n=1}^{N}$ from a large graph. These subgraphs will be incorporated into our prompt template and designed instructions to enable LLMs to generate the corresponding graph summaries $\{S_{n}\}_{n=1}^{N}$. In this study, we utilize open LLM, Qwen2-72B (Yang et al., 2024a), as the graph summary generator. Combining realistic unlabelled graph data, we generate and curate large-scale graph-summary pair data***https://github.com/ZhuYun97/GraphCLIP/summary/ which contains over 0.2B tokens across academic (He et al., 2023; Wang et al., 2020a; Yang et al., 2016), e-commerce (Hu et al., 2020), and social domains (Huang et al., 2024). For detailed instructions for graph summary generation, please refer to the Appendix C.
@@ 4.1. Datasets
Table 1. Statistics of Text-Attributed Graph datasets. $\mathcal{G}_{\text{source}}$ denotes source datasets, and $\mathcal{G}_{\text{target}}$ indicates target datasets.
@@ 4.3.1. Node Classification
For the LLM baselines, we directly apply them to the source data, as they have already been pretrained on extensive corpora, and we find that continuing to pretrain them on our source data deteriorates performance on the target data. For GraphGPT, we utilize their released checkpointhttps://github.com/HKUDS/GraphGPT and conduct experiments under our settings. For other methods, we use their official codes and pretrain models using the source datasets specified in Table 1 before applying them to the target datasets. For data splitting, we use the public split for the WikiCS dataset, while for other datasets, we randomly select 20% as testing samples. We report the mean accuracy along with the standard deviation after five runs with different random seeds.
@@
Table 3. Zero-shot inference for link prediction of different methods across target datasets.
@@ Appendix A The Dilemma of Vanilla Contrastive Loss
In a binary classification setting, let $(Z_{1},Z_{2})$ be normally distributed as $\mathcal{N}(0,I_{2})$. Assign $Y=1$ if $Z_{1}\geq 0$. For data augmentation, $Z_{2}$ is scaled by a normal distribution:
(12) $\tau_{\alpha}(Z)$ $=\left(Z_{1},\alpha\cdot Z_{2}\right)$
$\alpha$ $\sim\mathcal{N}(0,1)$
This creates a set of transformation-induced domains =$\mathcal{B}=\{\mathcal{G}_{m}:\mathcal{G}_{m}=(Z_{1},m\cdot Z_{2})\,|\,m\in\mathbb{R}\}$. For any $\zeta\geq 0$, there is a representation $g$ and two specific domains, $\mathcal{G}_{m}$ and $\mathcal{G}_{m^{\prime}}$, such that:
@@
However, $g$ demonstrates significantly varied performance across $\mathcal{G}_{m}$ and $\mathcal{G}_{m^{\prime}}$:
@@
where $\mathcal{L}_{\text{AL}}$ denotes alignment loss in the vanilla contrastive loss (Wang and Isola, 2020), $\mathcal{R}$ denotes supervised risk of binary classification. This demonstrates that even with a low alignment loss, the model’s performance can vary significantly across different augmented domains. This variability arises because a low $\mathcal{L}_{\text{AL}}$ is achieved through averaged alignments rather than uniform ones, causing representations in certain less frequently selected domains to experience substantial alignment losses.
Proof. For $\zeta\geq 0$, define $t=\sqrt{\zeta}/2$ and let $g(z_{1},z_{2})=z_{1}+tz_{2}.$ The alignment loss becomes:
(15) $\mathcal{L}_{\text{AL }}(g;\mathcal{G},\pi)=t^{2}\mathbb{E}Z_{2}^{2}\underset{\left(\alpha_{1},\alpha_{2}\right)\sim\mathcal{N}(0,1)^{2}}{\mathbb{E}}\left(\alpha_{1}-\alpha_{2}\right)^{2}=2t^{2}<\zeta.$
Setting $m=0$ and $m^{\prime}=1/t$, we have:
(16) $\mathcal{R}\left(g;\mathcal{G}_{m}\right)=0$
@@
(17) $\mathcal{R}\left(g;\mathcal{G}_{m^{\prime}}\right)=$
$P\left(Z_{1}<0,Z_{1}+Z_{2}\geq 0\right)+P\left(Z_{1}\geq 0,Z_{1}+Z_{2}\leq 0\right)=\frac{1}{4}$
@@ Appendix B Theoretical Analysis of Invariant Alignment Loss
When $\mathcal{L}_{\text{IAL}}$ is minimized to a small value, it signifies that $\mathcal{R}\left(c\circ g;\mathcal{G}\tau\right)$ remains consistent across various augmentation functions $\tau$, suggesting that the optimal representation for $\mathcal{G}_{\tau}$ closely resembles that of $\mathcal{G}_{\tau^{\prime}}$. In other words, representations with smaller $\mathcal{L}_{\text{IAL}}$ are more likely to yield similar linear optimal predictors across different domains, a characteristic that the original alignment loss lacks.
@@ Appendix D Datasets
WikiCS (Mernyei and Cangea, 2020) dataset is a Wikipedia-based dataset designed for benchmarking Graph Neural Networks, comprising 10 computer science branches as classes characterized by high connectivity. Node features are extracted from the corresponding article textsWe obtain the raw texts of each node from https://github.com/pmernyei/wiki-cs-dataset..
@@ G.1. Pretraining Phase
For GraphCLIP, only a few hyperparameters need to be adjusted. In our main experiments, we utilize the AdamW (Loshchilov and Hutter, [n. d.]) optimizer with both learning rate and weight decay set to 1e-5. The graph model employed is GraphGPS (Rampášek et al., 2022), consisting of 12 layers with a hidden size of 1024. For the text model, we use a fine-tuned version of MiniLM §§§https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 (Wang et al., 2020b), featuring 6 layers with a hidden size of 384. To align both models in a unified subspace, a projector is applied to transform the graph model’s 1024 dimensions to match the 384 dimensions of the text model. During pretraining, we optimize only the parameters of the graph model and projector, keeping the text model frozen to reduce training costs and mitigate catastrophic forgetting. Pretraining is conducted over 30 epochs with a batch size of 800 per GPU, utilizing eight A100-40G GPUs for pretraining within 7 hours. We will release our pretrained checkpoint https://github.com/ZhuYun97/GraphCLIP/checkpoints/ after the anonymous phase.
@@ G.2. Zero-shot Learning
In zero-shot learning, we incorporate label information into label-specific sentences to align with the pretraining format. Table 8 presents various prompts designed for different datasets. Placeholders are marked in violet font: “{class}” represents the label text for the target node, and “{class_desc}” is a descriptive sentence generated by LLMs to elaborate on the label. For detailed cases, please refer to our anonymous repository https://github.com/ZhuYun97/GraphCLIP/data/load.py.